# On the Enumeration of Generalized Cospectral Mates of Graphs

### v1

@@ Abstract
This paper investigates the enumeration of generalized cospectral mates of simple graphs, where the generalized spectrum consists of the spectra of a graph and its complement. Moving beyond the classical problem of identifying graphs determined by their generalized spectrum, we address the more quantitative question of how many non-isomorphic graphs can share the same generalized spectrum. Our approach is based on arithmetic constraints derived from the Smith Normal Form (SNF) of the walk matrix $W(G)$, which lead to a tight upper bound on the number of generalized cospectral mates of a graph. Our upper bound applies to a much broader class of graphs than those previously shown to have no generalized cospectral mates (determined by generalized spectrum). Consequently, this work extends the family of graphs for which strong and informative spectral uniqueness results are available.
@@ 1 Introduction
Wang and Xu [13] considered this problem in the context of generalized spectrum, defined as the pair $(\operatorname{Spec}(G),\operatorname{Spec}(\overline{G}))$. Two or more graphs are generalized cospectral if they have the same generalized spectrum. While isomorphic graphs are clearly generalized cospectral, the converse does not hold in general. Non-isomorphic graphs that are generalized cospectral are referred to as generalized cospectral mates. A graph is said to be determined by its generalized spectrum (DGS) if it has no generalized cospectral mate.
Recent advances have leveraged the properties of the walk matrix to address these questions. For a graph $G$ with $n$ vertices, the walk matrix $W(G)$ is the $n\times n$ matrix with columns $e,\,A(G)e,\,\dots,\,A(G)^{n-1}e$, where $e$ is the all-ones column vector of length $n$. A graph $G$ is controllable if $W(G)$ is non-singular. Wang [11] introduced a large family of controllable graphs and proved that all graphs in this family are DGS. Extending these ideas to oriented graphs, Qiu et al. [6] introduced a family of oriented graphs determined by their generalized skew spectra.
@@
Let $G\in\mathcal{F}_{n}$, and let $H_{1}$ and $H_{2}$ be graphs that are generalized cospectral with $G$. Let $Q_{1}$ and $Q_{2}$ be the rational orthogonal matrices satisfying $Q_{i}^{T}A(G)Q_{i}=A(H_{i})$ and $Q_{i}e=e$ for $i=1,2$. If $\ell(Q_{1})=\ell(Q_{2})$, then $H_{1}\cong H_{2}$.
@@ 2 Preliminaries
$\Gamma(G)=\{Q\in O_{n}(\mathbb{Q})\mid Q^{T}A(G)Q=A(H)\text{ for graph }H\},$
@@
[8, 14] Let $G$ be a graph, and let $H_{1}$ and $H_{2}$ be graphs generalized cospectral with $G$. Let $Q_{1},Q_{2}\in\Gamma(G)$ be the corresponding rational orthogonal matrices such that $Q_{i}^{T}A(G)Q_{i}=A(H_{i})$ and $Q_{i}e=e$ for $i=1,2$. If $Q_{2}=Q_{1}P$, where $P$ is a permutation matrix, then $H_{1}\cong H_{2}$.
@@
For any matrix $Q\in\Gamma(G)$, we have $Q^{T}W(G)=W(H)$ for some graph $H$ by Theorem 3. It follows that $\ell(Q)$ must divide $\operatorname{gcd}(d_{n}(W(G)),d_{n}(W(H)))$, and therefore it must divide $d_{n}(W(G))$. However, certain arithmetic properties of $W(G)$ can restrict which divisors of $d_{n}(W(G))$ are actually attainable as levels for matrices in $\Gamma(G)$.
@@ 3 Arithmetic Constraints on Matrices in $O_{n}(\mathbb{Q})$
where $y$ is any integer in $\mathbb{Z}/p^{k}\mathbb{Z}$ and $w^{\prime}=[0,\dots,0,1]^{T}$.
@@
    $u^{T}u\equiv v^{T}v\equiv 0\pmod{p^{2k}}$,
then $u^{T}v\equiv 0\pmod{p^{2k}}$.
@@
$(u-\alpha v)^{T}(u-\alpha v)\equiv 0\pmod{p^{2k}}$,
@@
$u^{T}u-2\alpha u^{T}v+\alpha^{2}v^{T}v\equiv 0\pmod{p^{2k}}$.
Since $u^{T}u\equiv 0\pmod{p^{2k}}$ and $v^{T}v\equiv 0\pmod{p^{2k}}$, we have
$2\alpha u^{T}v\equiv 0\pmod{p^{2k}}$.
As $p$ is odd and $\alpha$ is not divisible by $p$, it follows that $2\alpha$ is invertible in $\mathbb{Z}/p^{2k}\mathbb{Z}$. Hence $u^{T}v\equiv 0\pmod{p^{2k}}$.
@@
    $\operatorname{ord}_{2}(u^{T}u)=\operatorname{ord}_{2}(v^{T}v)=2k$,
@@
    $\operatorname{ord}_{2}(u^{T}e)=\operatorname{ord}_{2}(v^{T}e)=k$,
then $u^{T}v\equiv 0\pmod{2^{2k}}$.
@@
Based upon our assumptions, we have: (i) $u^{T}u=2^{2k}a$, (ii) $v^{T}v=2^{2k}b$. (iii) $u^{T}e=2^{k}c$, (iv) $v^{T}e=2^{k}d$, where $a$, $b$, $c$, and $d$ are odd integers. From condition (1), we have
@@
for some integer vector $m$. Multiplying each side by $e^{T}$ gives
$e^{T}u-\alpha e^{T}v=2^{k}e^{T}m.$
Substituting the values of $e^{T}u$ and $e^{T}v$ gives
$2^{k}c-\alpha(2^{k}d)=2^{k}e^{T}m,$
@@
$c-\alpha d=e^{T}m.$
Note that since $c$ and $\alpha d$ are both odd, their difference $c-\alpha d$ is even. Thus, $e^{T}m$ is even. Moreover, we can express $m^{T}m$ as
$m^{T}m$ $=\sum_{i=1}^{n}m_{i}^{2}$
@@
$\equiv e^{T}m\pmod{2}$,
where $m_{i}$ is the $i^{\text{th}}$ entry of $m$. Since $e^{T}m$ is even, it follows that $m^{T}m$ is also even.
@@
$(u-\alpha v)^{T}(u-\alpha v)$ $=(2^{k}m)^{T}(2^{k}m)$
$u^{T}u-2\alpha u^{T}v+\alpha^{2}v^{T}v$ $=2^{2k}m^{T}m.$
Substituting the values of $u^{T}u$ and $v^{T}v$ gives:
$2^{2k}a-2\alpha u^{T}v+\alpha^{2}(2^{2k}b)$ $=2^{2k}m^{T}m$
$-2\alpha u^{T}v$ $=2^{2k}(m^{T}m-(a+\alpha^{2}b))$
$-\alpha u^{T}v$ $=2^{2k-1}(m^{T}m-(a+\alpha^{2}b)).$
Note that since $a$ and $\alpha^{2}b$ are both odd, their sum $a+\alpha^{2}b$ is even. Since $m^{T}m$ is also even, it follows that $m^{T}m-(a+\alpha^{2}b)$ is even. Thus, we conclude
$\alpha u^{T}v\equiv 0\pmod{2^{2k}}$.
Since $\alpha$ is odd, it is invertible in $\mathbb{Z}/2^{2k}\mathbb{Z}$. Thus, $u^{T}v\equiv 0\pmod{2^{2k}}$. ∎
@@
Let $\ell=\ell(Q)$. Since $Q$ is orthogonal, $Q^{T}Q=I$, and therefore $\bar{Q}^{T}\bar{Q}=(\ell Q)^{T}(\ell Q)=\ell^{2}I.$ As the $(i,j)$-th entry of $\bar{Q}^{T}\bar{Q}$ is $v_{i}^{T}v_{j}$, it follows that
$v_{i}^{T}v_{j}=\begin{cases}\ell^{2}&\text{if }i=j,\\ 0&\text{if }i\neq j.\end{cases}$
Since $k=\operatorname{ord}_{p}(\ell)$, we have $p^{k}\parallel\ell$ and thus $p^{2k}\parallel\ell^{2}$. It follows that $v_{i}^{T}v_{j}\equiv 0\pmod{p^{2k}}$ for all $i,j$ while $p^{2k}\parallel v_{i}^{T}v_{i}$ for all $i$. This proves (1) and (2).
To prove (3), since $Q$ is regular and orthogonal, we have $Q^{T}e=e$. Multiplying by $\ell$ gives $\bar{Q}^{T}e=\ell e.$ The $i^{\text{th}}$ component of $\bar{Q}^{T}e$ is $v_{i}^{T}e=\ell$. Since $p^{k}\parallel\ell$, we conclude that $p^{k}\parallel v_{i}^{T}e$. ∎
@@
Let $R=\bar{Q}_{1}^{T}\bar{Q}_{2}$ be an integral matrix with entries $r_{ij}=u_{i}^{T}v_{j}$, where $u_{i}$ and $v_{j}$ denote the $i^{\text{th}}$ and $j^{\text{th}}$ columns of $\bar{Q}_{1}$ and $\bar{Q}_{2}$, respectively. We show that $r_{ij}\equiv 0\pmod{p^{2k}}$ for all $i,j$.
By the definition of level, there exist column vectors $u^{\prime}$ of $\bar{Q}_{1}$ and $v^{\prime}$ of $\bar{Q}_{2}$ such that $u^{\prime}\not\equiv 0\pmod{p}$ and $v^{\prime}\not\equiv 0\pmod{p}$. Since $u^{\prime}$ and $v^{\prime}$ are both scalar multiples of $w$ over $\mathbb{Z}/p^{k}\mathbb{Z}$, there exists an integer $\alpha$ not divisible by $p$ and an integer vector $z$ such that:
$u^{\prime}\equiv\alpha v^{\prime}+p^{k}z\pmod{p^{2k}}$, (5)
@@
$u_{i}\equiv c_{i}u^{\prime}+p^{k}x_{i}\pmod{p^{2k}}$,
$v_{j}\equiv d_{j}v^{\prime}+p^{k}y_{j}\pmod{p^{2k}}$,
@@
Since $Q_{1}$ and $Q_{2}$ satisfy the conditions of Lemma 13, we have that $\operatorname{ord}_{p}(u^{\prime T}u^{\prime})=\operatorname{ord}_{p}(v^{\prime T}v^{\prime})=2k$ and $\operatorname{ord}_{p}(u^{\prime T}e)=\operatorname{ord}_{p}(v^{\prime T}e)=k$. Therefore, we can apply Lemma 11 (for odd primes) or Lemma 12 (for $p=2$) to establish that $u^{\prime T}v^{\prime}\equiv 0\pmod{p^{2k}}$. Multiplying Eq. (5) by $u^{\prime T}$ yields:
$u^{\prime T}u^{\prime}$ $\equiv u^{\prime T}(\alpha v^{\prime}+p^{k}z)$
$\equiv\alpha u^{\prime T}v^{\prime}+p^{k}u^{\prime T}z\pmod{p^{2k}}$.
Since $u^{\prime T}u^{\prime}\equiv 0\pmod{p^{2k}}$ and $u^{\prime T}v^{\prime}\equiv 0\pmod{p^{2k}}$, it follows that $p^{k}u^{\prime T}z\equiv 0\pmod{p^{2k}}$, which implies $u^{\prime T}z\equiv 0\pmod{p^{k}}$.
Next, we examine the product $u_{i}^{T}u^{\prime}$. Substituting the expression for $u_{i}$ and $u^{\prime}$ gives:
$u_{i}^{T}u^{\prime}$ $\equiv(c_{i}u^{\prime}+p^{k}x_{i})^{T}(\alpha v^{\prime}+p^{k}z)$
$\equiv c_{i}\alpha u^{\prime T}v^{\prime}+c_{i}p^{k}u^{\prime T}z+\alpha p^{k}x_{i}^{T}v^{\prime}+p^{2k}x_{i}^{T}z\pmod{p^{2k}}$.
Note that $u^{\prime T}v^{\prime}\equiv 0\pmod{p^{2k}}$. Since $u^{\prime T}z\equiv 0\pmod{p^{k}}$, the term $c_{i}p^{k}u^{\prime T}z$ is divisible by $p^{2k}$. Furthermore, from Lemma 13, we know $u_{i}^{T}u^{\prime}\equiv 0\pmod{p^{2k}}$. Thus, we get:
$0\equiv\alpha p^{k}x_{i}^{T}v^{\prime}\pmod{p^{2k}}$.
Since $\alpha\not\equiv 0\pmod{p}$, we conclude $x_{i}^{T}v^{\prime}\equiv 0\pmod{p^{k}}$. By a similar argument, we can prove that $y_{j}^{T}u^{\prime}\equiv 0\pmod{p^{k}}$. Finally, for any columns $u_{i}$ and $v_{j}$, we have:
$u_{i}^{T}v_{j}$ $\equiv(c_{i}u^{\prime}+p^{k}x_{i})^{T}(d_{j}v^{\prime}+p^{k}y_{j})$
$\equiv c_{i}d_{j}u^{\prime T}v^{\prime}+c_{i}p^{k}u^{\prime T}y_{j}+d_{j}p^{k}x_{i}^{T}v^{\prime}+p^{2k}x_{i}^{T}y_{j}\pmod{p^{2k}}$.
We established that $u^{\prime T}y_{j}\equiv 0\pmod{p^{k}}$ and $x_{i}^{T}v^{\prime}\equiv 0\pmod{p^{k}}$. Since $u^{\prime T}v^{\prime}\equiv 0\pmod{p^{2k}}$, every term on the right-hand side vanishes modulo $p^{2k}$. Thus, we conclude that $u_{i}^{T}v_{j}\equiv 0\pmod{p^{2k}}$ for all $i,j$, which implies $R\equiv 0\pmod{p^{2k}}$. Substituting the definition of $R$, we have
$\ell(Q_{1})\ell(Q_{2})Q_{1}^{T}Q_{2}\equiv 0\pmod{p^{2k}}$.
As $Q_{1}^{T}Q_{2}$ is a rational matrix, by Lemma 5, it follows that $\ell(Q_{1}^{T}Q_{2})\mid\frac{\ell(Q_{1})\ell(Q_{2})}{p^{2k}}$. Since $p^{k}\parallel\ell(Q_{1})$ and $p^{k}\parallel\ell(Q_{2})$, it follows that $p^{2k}\parallel\ell(Q_{1})\ell(Q_{2})$. Consequently, $p\nmid\frac{\ell(Q_{1})\ell(Q_{2})}{p^{2k}}$, which implies $p\nmid\ell(Q_{1}^{T}Q_{2})$, completing the proof. ∎
@@
be the prime factorization of $\ell$. Suppose that for each $j\in\{1,\dots,m\}$, there exists a vector $w_{j}\not\equiv 0\pmod{p_{j}}$ such that columns of both $\bar{Q}_{1}$ and $\bar{Q}_{2}$ is a scalar multiple of $w_{j}$ over the ring $\mathbb{Z}/p_{j}^{k_{j}}\mathbb{Z}$. Then $Q_{1}^{T}Q_{2}$ is a permutation matrix.

### v2

# On the number of generalized cospectral mates of graphs

@@ Abstract
This paper establishes an upper bound on the number of generalized cospectral mates of simple graphs, where the generalized spectrum consists of the spectrum of a graph and its complement. Moving beyond the classical problem of identifying graphs determined by their generalized spectrum, we address the more quantitative question of how many non-isomorphic graphs can share the same generalized spectrum. Our approach is based on arithmetic constraints derived from the Smith Normal Form (SNF) of the walk matrix, which leads to a tight upper bound on the number of generalized cospectral mates of a graph. Our upper bound applies to a much broader class of graphs than those previously shown to have no generalized cospectral mates (graphs determined by generalized spectrum). Consequently, this work extends the family of graphs for which strong and informative spectral uniqueness results are available.
@@ 1 Introduction
Wang and Xu [11] considered this problem in the context of generalized spectrum, defined as the pair $(\operatorname{Spec}(G),\operatorname{Spec}(\overline{G}))$. Two or more graphs are generalized cospectral if they have the same generalized spectrum. While isomorphic graphs are clearly generalized cospectral, the converse does not hold in general. Pairwise non-isomorphic graphs that are generalized cospectral are referred to as generalized cospectral mates. A graph is said to be determined by its generalized spectrum (DGS) if it has no generalized cospectral mate.
Recent advances have leveraged the properties of the walk matrix to address these questions. For a graph $G$ with $n$ vertices, the walk matrix $W(G)$ is the $n\times n$ matrixa with columns $e,\,A(G)e,\,\dots,\,A(G)^{n-1}e$, where $e$ is the all-ones column vector of length $n$. A graph $G$ is controllable if $W(G)$ is non-singular. Wang [12] introduced a large family of controllable graphs and proved that all graphs in this family are DGS. Extending these ideas to oriented graphs, Qiu et al. [6] introduced a family of oriented graphs determined by their generalized skew spectra.
@@
Let $G\in\mathcal{F}_{n}$, and let $H_{1}$ and $H_{2}$ be graphs that are generalized cospectral with $G$. Let $Q_{1}$ and $Q_{2}$ be the rational orthogonal matrices satisfying $Q_{i}^{\mathrm{T}}A(G)Q_{i}=A(H_{i})$ and $Q_{i}e=e$ for $i=1,2$. If $\ell(Q_{1})=\ell(Q_{2})$, then $H_{1}\cong H_{2}$.
@@ 2 Preliminaries
$\Gamma(G)=\{Q\in O_{n}(\mathbb{Q})\mid Q^{\mathrm{T}}A(G)Q=A(H)\text{ for graph }H\},$
@@
[8, 13] Let $G$ be a graph, and let $H_{1}$ and $H_{2}$ be graphs generalized cospectral with $G$. Let $Q_{1},Q_{2}\in\Gamma(G)$ be the corresponding rational orthogonal matrices such that $Q_{i}^{\mathrm{T}}A(G)Q_{i}=A(H_{i})$ and $Q_{i}e=e$ for $i=1,2$. If $Q_{2}=Q_{1}P$, where $P$ is a permutation matrix, then $H_{1}\cong H_{2}$.
@@
For any matrix $Q\in\Gamma(G)$, we have $Q^{\mathrm{T}}W(G)=W(H)$ for some graph $H$ by Theorem 3. It follows that $\ell(Q)$ must divide $\operatorname{gcd}(d_{n}(W(G)),d_{n}(W(H)))$, and therefore it must divide $d_{n}(W(G))$. However, certain arithmetic properties of $W(G)$ can restrict which divisors of $d_{n}(W(G))$ are actually attainable as levels for matrices in $\Gamma(G)$.
@@ 3 Arithmetic Constraints on Matrices in $O_{n}(\mathbb{Q})$
where $y$ is any integer in $\mathbb{Z}/p^{k}\mathbb{Z}$ and $w^{\prime}=[0,\dots,0,1]^{\mathrm{T}}$.
@@
    $u^{\mathrm{T}}u\equiv v^{\mathrm{T}}v\equiv 0\pmod{p^{2k}}$,
then $u^{\mathrm{T}}v\equiv 0\pmod{p^{2k}}$.
@@
$(u-\alpha v)^{\mathrm{T}}(u-\alpha v)\equiv 0\pmod{p^{2k}}$,
@@
$u^{\mathrm{T}}u-2\alpha u^{\mathrm{T}}v+\alpha^{2}v^{\mathrm{T}}v\equiv 0\pmod{p^{2k}}$.
Since $u^{\mathrm{T}}u\equiv 0\pmod{p^{2k}}$ and $v^{\mathrm{T}}v\equiv 0\pmod{p^{2k}}$, we have
$2\alpha u^{\mathrm{T}}v\equiv 0\pmod{p^{2k}}$.
As $p$ is odd and $\alpha$ is not divisible by $p$, it follows that $2\alpha$ is invertible in $\mathbb{Z}/p^{2k}\mathbb{Z}$. Hence $u^{\mathrm{T}}v\equiv 0\pmod{p^{2k}}$.
@@
    $\operatorname{ord}_{2}(u^{\mathrm{T}}u)=\operatorname{ord}_{2}(v^{\mathrm{T}}v)=2k$,
@@
    $\operatorname{ord}_{2}(u^{\mathrm{T}}e)=\operatorname{ord}_{2}(v^{\mathrm{T}}e)=k$,
then $u^{\mathrm{T}}v\equiv 0\pmod{2^{2k}}$.
@@
Based upon our assumptions, we have: (i) $u^{\mathrm{T}}u=2^{2k}a$, (ii) $v^{\mathrm{T}}v=2^{2k}b$. (iii) $u^{\mathrm{T}}e=2^{k}c$, (iv) $v^{\mathrm{T}}e=2^{k}d$, where $a$, $b$, $c$, and $d$ are odd integers. From condition (1), we have
@@
for some integer vector $m$. Multiplying each side by $e^{\mathrm{T}}$ gives
$e^{\mathrm{T}}u-\alpha e^{\mathrm{T}}v=2^{k}e^{\mathrm{T}}m.$
Substituting the values of $e^{\mathrm{T}}u$ and $e^{\mathrm{T}}v$ gives
$2^{k}c-\alpha(2^{k}d)=2^{k}e^{\mathrm{T}}m,$
@@
$c-\alpha d=e^{\mathrm{T}}m.$
Note that since $c$ and $\alpha d$ are both odd, their difference $c-\alpha d$ is even. Thus, $e^{\mathrm{T}}m$ is even. Moreover, we can express $m^{\mathrm{T}}m$ as
$m^{\mathrm{T}}m$ $=\sum_{i=1}^{n}m_{i}^{2}$
@@
$\equiv e^{\mathrm{T}}m\pmod{2}$,
where $m_{i}$ is the $i^{\text{th}}$ entry of $m$. Since $e^{\mathrm{T}}m$ is even, it follows that $m^{\mathrm{T}}m$ is also even.
@@
$(u-\alpha v)^{\mathrm{T}}(u-\alpha v)$ $=(2^{k}m)^{\mathrm{T}}(2^{k}m)$
$u^{\mathrm{T}}u-2\alpha u^{\mathrm{T}}v+\alpha^{2}v^{\mathrm{T}}v$ $=2^{2k}m^{\mathrm{T}}m.$
Substituting the values of $u^{\mathrm{T}}u$ and $v^{\mathrm{T}}v$ gives:
$2^{2k}a-2\alpha u^{\mathrm{T}}v+\alpha^{2}(2^{2k}b)$ $=2^{2k}m^{\mathrm{T}}m$
$-2\alpha u^{\mathrm{T}}v$ $=2^{2k}(m^{\mathrm{T}}m-(a+\alpha^{2}b))$
$-\alpha u^{\mathrm{T}}v$ $=2^{2k-1}(m^{\mathrm{T}}m-(a+\alpha^{2}b)).$
Note that since $a$ and $\alpha^{2}b$ are both odd, their sum $a+\alpha^{2}b$ is even. Since $m^{\mathrm{T}}m$ is also even, it follows that $m^{\mathrm{T}}m-(a+\alpha^{2}b)$ is even. Thus, we conclude
$\alpha u^{\mathrm{T}}v\equiv 0\pmod{2^{2k}}$.
Since $\alpha$ is odd, it is invertible in $\mathbb{Z}/2^{2k}\mathbb{Z}$. Thus, $u^{\mathrm{T}}v\equiv 0\pmod{2^{2k}}$. ∎
@@
Let $\ell=\ell(Q)$. Since $Q$ is orthogonal, $Q^{\mathrm{T}}Q=I$, and therefore $\bar{Q}^{\mathrm{T}}\bar{Q}=(\ell Q)^{\mathrm{T}}(\ell Q)=\ell^{2}I.$ As the $(i,j)$-th entry of $\bar{Q}^{\mathrm{T}}\bar{Q}$ is $v_{i}^{\mathrm{T}}v_{j}$, it follows that
$v_{i}^{\mathrm{T}}v_{j}=\begin{cases}\ell^{2}&\text{if }i=j,\\ 0&\text{if }i\neq j.\end{cases}$
Since $k=\operatorname{ord}_{p}(\ell)$, we have $p^{k}\parallel\ell$ and thus $p^{2k}\parallel\ell^{2}$. It follows that $v_{i}^{\mathrm{T}}v_{j}\equiv 0\pmod{p^{2k}}$ for all $i,j$ while $p^{2k}\parallel v_{i}^{\mathrm{T}}v_{i}$ for all $i$. This proves (1) and (2).
To prove (3), since $Q$ is regular and orthogonal, we have $Q^{\mathrm{T}}e=e$. Multiplying by $\ell$ gives $\bar{Q}^{\mathrm{T}}e=\ell e.$ The $i^{\text{th}}$ component of $\bar{Q}^{\mathrm{T}}e$ is $v_{i}^{\mathrm{T}}e=\ell$. Since $p^{k}\parallel\ell$, we conclude that $p^{k}\parallel v_{i}^{\mathrm{T}}e$. ∎
@@
Let $R=\bar{Q}_{1}^{\mathrm{T}}\bar{Q}_{2}$ be an integral matrix with entries $r_{ij}=u_{i}^{\mathrm{T}}v_{j}$, where $u_{i}$ and $v_{j}$ denote the $i^{\text{th}}$ and $j^{\text{th}}$ columns of $\bar{Q}_{1}$ and $\bar{Q}_{2}$, respectively. We show that $r_{ij}\equiv 0\pmod{p^{2k}}$ for all $i,j$.
By the definition of level, there exist column vectors $u^{\prime}$ of $\bar{Q}_{1}$ and $v^{\prime}$ of $\bar{Q}_{2}$ such that $u^{\prime}\not\equiv 0\pmod{p}$ and $v^{\prime}\not\equiv 0\pmod{p}$. Since both $u^{\prime}$ and $v^{\prime}$ are also scalar multiples of $w$ over $\mathbb{Z}/p^{k}\mathbb{Z}$, there exists an integer $\alpha\not\equiv 0\pmod{p}$ such that:
$u^{\prime}\equiv\alpha v^{\prime}\pmod{p^{k}}$, (5)
@@
$u_{i}\equiv c_{i}u^{\prime}+p^{k}x_{i}\pmod{p^{2k}}$, (6)
$v_{j}\equiv d_{j}v^{\prime}+p^{k}y_{j}\pmod{p^{2k}}$, (7)
@@
Since $Q_{1}$ and $Q_{2}$ satisfy the conditions of Lemma 13, it follows that $u_{i}^{\mathrm{T}}u^{\prime}\equiv 0\pmod{p^{2k}}$ and $v_{j}^{\mathrm{T}}v^{\prime}\equiv 0\pmod{p^{2k}}$ for each column $u_{i}$ and $v_{j}$. Moreover, $\operatorname{ord}_{p}({u^{\prime}}^{\mathrm{T}}u^{\prime})=\operatorname{ord}_{p}({v^{\prime}}^{\mathrm{T}}v^{\prime})=2k$ and $\operatorname{ord}_{p}({u^{\prime}}^{\mathrm{T}}e)=\operatorname{ord}_{p}({v^{\prime}}^{\mathrm{T}}e)=k$. Therefore, we can apply Lemma 11 (for odd primes) or Lemma 12 (for $p=2$) to establish that ${u^{\prime}}^{\mathrm{T}}v^{\prime}\equiv 0\pmod{p^{2k}}$.
Multiplying Eq. (6) by ${u^{\prime}}^{\mathrm{T}}$ yields:
${u^{\prime}}^{\mathrm{T}}u_{i}$ $\equiv{u^{\prime}}^{\mathrm{T}}(c_{i}u^{\prime}+p^{k}x_{i})\pmod{p^{2k}}$
$\equiv c_{i}{u^{\prime}}^{\mathrm{T}}u^{\prime}+p^{k}{u^{\prime}}^{\mathrm{T}}x_{i}\pmod{p^{2k}}$
As ${u^{\prime}}^{\mathrm{T}}u_{i}\equiv 0\pmod{p^{2k}}$ and $c_{i}{u^{\prime}}^{\mathrm{T}}u^{\prime}\equiv 0\pmod{p^{2k}}$, it follows that $p^{k}{u^{\prime}}^{\mathrm{T}}x_{i}\equiv 0\pmod{p^{2k}}$, which implies ${u^{\prime}}^{\mathrm{T}}x_{i}\equiv 0\pmod{p^{k}}$. Using Eq. (5), we get ${u^{\prime}}^{\mathrm{T}}x_{i}\equiv\alpha{v^{\prime}}^{\mathrm{T}}x_{i}\pmod{p^{k}}$. As $\alpha\not\equiv 0\pmod{p}$, we conclude that ${v^{\prime}}^{\mathrm{T}}x_{i}\equiv 0\pmod{p^{k}}$. Using a similar approach, we can prove that ${u^{\prime}}^{\mathrm{T}}y_{j}\equiv 0\pmod{p^{k}}$ for each $y_{j}$.
Finally, for any columns $u_{i}$ and $v_{j}$, we have:
$u_{i}^{\mathrm{T}}v_{j}$ $\equiv(c_{i}u^{\prime}+p^{k}x_{i})^{\mathrm{T}}(d_{j}v^{\prime}+p^{k}y_{j})$
$\equiv c_{i}d_{j}{u^{\prime}}^{\mathrm{T}}v^{\prime}+c_{i}p^{k}{u^{\prime}}^{\mathrm{T}}y_{j}+d_{j}p^{k}x_{i}^{\mathrm{T}}v^{\prime}+p^{2k}x_{i}^{\mathrm{T}}y_{j}\pmod{p^{2k}}$.
We established that ${u^{\prime}}^{\mathrm{T}}y_{j}\equiv 0\pmod{p^{k}}$ and ${v^{\prime}}^{\mathrm{T}}x_{i}\equiv 0\pmod{p^{k}}$. Since ${u^{\prime}}^{\mathrm{T}}v^{\prime}\equiv 0\pmod{p^{2k}}$, every term on the right-hand side vanishes modulo $p^{2k}$. Thus, we conclude that $u_{i}^{\mathrm{T}}v_{j}\equiv 0\pmod{p^{2k}}$ for all $i,j$, which implies $R\equiv 0\pmod{p^{2k}}$. Substituting the definition of $R$, we have
$\ell(Q_{1})\ell(Q_{2})Q_{1}^{\mathrm{T}}Q_{2}\equiv 0\pmod{p^{2k}}$.
As $Q_{1}^{\mathrm{T}}Q_{2}$ is a rational matrix, by Lemma 5, it follows that $\ell(Q_{1}^{\mathrm{T}}Q_{2})\mid\frac{\ell(Q_{1})\ell(Q_{2})}{p^{2k}}$. Since $p^{k}\parallel\ell(Q_{1})$ and $p^{k}\parallel\ell(Q_{2})$, it follows that $p^{2k}\parallel\ell(Q_{1})\ell(Q_{2})$. Consequently, $p\nmid\frac{\ell(Q_{1})\ell(Q_{2})}{p^{2k}}$, which implies $p\nmid\ell(Q_{1}^{\mathrm{T}}Q_{2})$, completing the proof. ∎
@@
be the prime factorization of $\ell$. Suppose that for each $j\in\{1,\dots,m\}$, there exists a vector $w_{j}\not\equiv 0\pmod{p_{j}}$ such that columns of both $\bar{Q}_{1}$ and $\bar{Q}_{2}$ is a scalar multiple of $w_{j}$ over the ring $\mathbb{Z}/p_{j}^{k_{j}}\mathbb{Z}$. Then $Q_{1}^{\mathrm{T}}Q_{2}$ is a permutation matrix.